# An ETH-Tight FPT Algorithm for Rejection-Proof Set Packing with Applications to Kidney Exchange

### v1

@@ Abstract
We exploit the sunflower lemma on a set packing formulation of the problem to give a kernel for this $\Sigma_{2}^{P}$-complete problem that is polynomial in $k$ for all constant values of $d$. We also provide a $2^{\mathcal{O}(k\log k)}+n^{\mathcal{O}(1)}$ algorithm based on it and show that this FPT algorithm is asymptotically optimal under the ETH. Further, we generalize the problem by including an additional positive integer $c$ in the input that naturally captures how much agents can modify a given cycle packing to reject it. For every constant $c$, the resulting problem simplifies from being $\Sigma_{2}^{P}$-complete to NP-complete. With a single-exponential algorithm for the setting where $c=1$, we show this to be strictly easier under the ETH than when $c=2$. In turn, we show that any $c\geq 2$ yields a problem that is essentially as hard as the original problem with $c$ unbounded. This displays an interesting discrepancy between the classical and parameterized complexity of the problem and gives a good view of what makes it hard.
@@ 1 Introduction
For patients with severely reduced kidney function, transplantation is often the preferred treatment [Burra07]. As it is possible to live with only one healthy kidney, live kidney donation is employed to reduce waiting times for sparsely available donor organs. In this setting, a willing donor offers one of their kidneys for transplantation to a patient. The patient and donor must be medically compatible for a successful transplantation. When the patient and donor are not compatible, they can opt to participate in a kidney exchange program (KEP).
Such programs aim to link multiple patient-donor pairs so that a donor from one pair is compatible with a patient from another pair. Their goal is to group sets of patient-donor pairs, such that within each group, a cyclic order of compatibility exists. Performing this cycle of transplants means that donors only undergo surgery if their corresponding patient receives a kidney in that exchange. Individual programs each have their own additional restrictions [Biro21]. It is common to restrict the length of transplantation cycles to some constant $d$, due to practical considerations which include doing the associated surgeries simultaneously.
We study these KEPs from an algorithmic perspective. Given a set of participating patient-donor pairs and the medical compatibilities between them, the algorithmic task becomes to find a combination of transplantation cycles that maximizes the total number of patients receiving a healthy kidney. We can model each patient-donor pair as a vertex and the compatibility of the donor of a pair $a$ to the patient of a pair $b$ as a directed edge from $a$ to $b$. We call a cycle of length at most dddd-cycle and use this notation to obtain the following algorithmic graph problem that describes KEPs in one of their most basic forms.
@@
This problem has already been studied from the perspective of parameterized algorithms in the literature, with fixed-parameter tractable (FPT) algorithms w.r.t. the input parameter $k$ and the number of neighborhood equivalence classes as examples [Hebert-JohnsonL24, MaitiD22].
@@
Extensions of the standard kidney exchange problem have also been studied. In this work, we consider one such extension which we now motivate and introduce. It is known that the fraction of patients to receive a healthy kidney can increase significantly with an increase in the total number of participating pairs [Agarwal19]. This motivates collaborations between multiple programs, from now on referred to as agents. Such collaborations, however, can be hindered by a mismatch between the individual interests of agents, as these need not align with the collective goal of helping as many patients as possible overall. In particular, a socially optimal solution—in which as many patients as possible are treated overall—is not guaranteed to help as many people from each individual agent as that agent could have achieved on their own. This may stop agents from participating in collaborative programs or, when they do participate, move them to withhold information from the collective.
Several ways of incorporating fairness into multi-agent KEPs have been studied [Agarwal19, Ashlagi15, Biro20, Carvalho23, Carvalho16, Hajaj15, Klimentova21]. Recently, Blom, Smeulders, and Spieksma [Blom24] also introduced a framework to capture fairness via a game-theoretical concept of rejections. They translate this into an algorithmic task that asks for a solution that is not rejected by any agent. We study the resulting algorithmic task. To introduce it, we first define the concept of rejections.
@@
This definition captures that the rejecting agent $i$ could modify a given cycle packing $\mathcal{C}$ into another packing $\mathcal{C}^{\prime}$ that benefits more of their own patients by only making changes pertaining to patients within their control. We call a packing that is not rejected by any agent rejection-proof. Then, a guarantee to always find such a solution also serves as a guarantee for agents that participation in the joint program is never detrimental to their personal interests. This led Blom et al. [Blom24] to introduce the following problem, whose parameterized complexity we investigate in this work.
@@
Rejection-Proof Kidney Exchange with $d$-cycles (RPKE-$d$) A directed graph $G$ where $V(G)=V_{1}\cup\ldots\cup V_{p}$ is partitioned among $p$ agents, and an interger $k$. Determine whether there is a rejection-proof set $\mathcal{C}$ of pairwise vertex-disjoint $d$-cycles such that $\mathcal{C}$ covers at least $k$ vertices. For ease of notation, we refer to any packing of pairwise vertex-disjoint $d$-cycles that covers at least $k$ vertices in an RPKE-$d$ instance as a candidate solution. Hence, a packing of $d$-cycles forms a solution to an instance if it is both a candidate solution and rejection-proof.
The above problem takes a leap in complexity compared to the single-agent problem KE-$d$. Whereas KE-$d$ is known to be NP-complete for all finite $d\geq 3$ [Abraham07], Blom et al. have shown RPKE-$d$ to be $\Sigma_{2}^{P}$-complete for $d\geq 3$, even when there are only two participating agents [Blom24, Theorem 1]. In fact, it can be shown that checking whether a given agent rejects a given candidate solution is already NP-hard since this is at least as hard as checking whether a given packing is optimal in the NP-hard setting of only one agent. Partly motivated by this observation, we also consider a more restricted definition of rejections.
For some integer $c\geq 0$, we say that an agent $c$-rejects a packing $\mathcal{C}$ if the rejection criterion from 1.1 can be satisfied with $|\mathcal{C}_{\mathrm{rej}}|\leq c$. We call a packing that no agent $c$-rejects $c$-rejection-proof. Now, we define the problem $c$-Rejection-Proof Kidney Exchange with $d$-cycles ($c$-RPKE-$d$) as the variant of the above RPKE-$d$ problem in which we aim to determine whether the input contains a candidate solution that is $c$-rejection-proof.
@@
We study the complexity of RPKE-$d$ and $c$-RPKE-$d$ parameterized by $k$, the number of vertices to be covered, and we present both positive and negative results. We prove our positive results on related set packing problems that generalize the two graph problems. This allows us to use some of the tools developed for such problems.
@@
We give an algorithm that, for constant values of $d$, solves RPKE-$d$ in $2^{\mathcal{O}(k\log k)}+n^{\mathcal{O}(1)}$ time on $n$-vertex graphs. It consists of two parts, the first of which is a kernelization algorithm. It takes an arbitrary instance as input and outputs an equivalent instance of size $\mathcal{O}\left((k^{d+1})^{d}\right)$ in polynomial time, via a careful agent-aware application of the sunflower lemma [Erdos60] to eliminate transplantation cycles. In set packing problems, removing sets during preprocessing typically comes at the risk of removing sets that are required to make an optimal solution, thereby possibly transforming YES-instances into NO-instances. The sunflower lemma provides a tool to circumvent this. Our rejection-proof problem settings introduce an additional type of behavior that needs care: transplantation cycles that do not belong to optimal solutions may still play a role in rejections. Removing the wrong cycles may cause an agent not to reject a certain candidate solution that it otherwise would have. We present a two-step reduction approach guaranteed to preserve the answer. The second part of the algorithm consists of brute-force on the reduced instance.
@@
We give a reduction from Subgraph Isomorphism to show that, under the Exponential Time Hypothesis (ETH), the algorithm is asymptotically optimal in terms of $n$ and $k$. We even show the stronger bound (independent of $k$) that a $2^{o(n\log n)}$ time algorithm for RPKE-$d$ or $c$-RPKE-$d$ would violate the ETH for any $c\geq 2$ and $d\geq 3$. We exploit the mechanism of rejections to facilitate our reduction from the Subgraph Isomorphism problem, for which such a superexponential ETH-based lower bound is known [CyganFGKMPS16]. This gives the rejection-proof kidney exchange problems a similar complexity status to other problems in the literature, such as Metric Dimension and Geodetic Set parameterized by vertex cover number, for which a brute-force algorithm on a polynomial kernel is also ETH-optimal [FoucaudGK0IST25].
@@
To complement the ETH-hardness of $c$-RPKE-$d$ with $c\geq 2$, we give a $3^{n}\cdot n^{\mathcal{O}(1)}$ time algorithm for the problem when $c=1$. This allows us to see surprising jumps in the algorithmic complexity of $c$-RPKE-$d$ for varying values of $c$. While no big difference in complexity is obtained between $c=1$ and $c=0$ (in which case the problem can be solved in $2^{n}\cdot n^{\mathcal{O}(1)}$ time with a standard application of dynamic programming over subsets [CyganFGKMPS16, Chapter 6]), our results show that the problem does become computationally harder from a parameterized perspective when $c=2$ (assuming the ETH). We also see that no more jumps in parameterized complexity appear for larger values of $c$. The surprising aspect here is that this one discrete jump in the parameterized complexity of $c$-RPKE-$d$ does not coincide with the jump in the classical complexity of the problem between being contained in NP vs. not being contained in NP: we have seen that, for constant $d\geq 3$, the $c$-RPKE-$d$ problem is contained in NP for every constant value of $c$, while the RPKE-$d$ problem is $\Sigma_{2}^{P}$-complete.
@@
The remainder of this work is organized as follows. We start with some preliminaries in section 2. Next, we give a polynomial kernel and an FPT algorithm for a set packing generalization of RPKE-$d$ in LABEL:sec:kernel, and we present a matching ETH-lower bound in LABEL:sec:eth-lower-bound. In LABEL:sec:1-rejection-algo, we give a single-exponential algorithm under $1$-rejections and we conclude in LABEL:sec:conclusion. The proofs of statements marked $(\bigstar)$ are deferred to the appendix.
@@ 2 Preliminaries
All graphs considered are directed, unless stated otherwise. We use standard notation for graphs and concepts from parameterized complexity; see the textbook by Cygan et al. [CyganFKLMPPS15]. We write $\langle v_{1},\ldots,v_{\ell}\rangle$ to denote the directed cycle with edge set $\{(v_{1},v_{2}),\ldots,(v_{\ell-1},v_{\ell}),(v_{\ell},v_{1})\}$.
@@
Next, we present the sunflower lemma by Erdős and Rado [Erdos60] and a related definition of sunflowers in set systems. In a set system $(U,\mathcal{S})$, a sunflower $\mathcal{F}$ with $z$ petals and core $Y$ is a collection of sets $S_{1},\ldots,S_{z}\in\mathcal{S}$ such that $S_{i}\cap S_{j}=Y$ for all $i\neq j$. We refer to a set $S_{i}\setminus Y$ as the petal corresponding to set $S_{i}$ and we require all petals to be non-empty. We state a variant of it as proven in the textbook by Flum and Grohe [FlumGrohe12, page 212].

### v2

@@ Abstract
We exploit the sunflower lemma on a set packing formulation of the problem to give a kernel for this $\Sigma_{2}^{P}$-complete problem that is polynomial in $k$ for all constant values of $d$. We also provide a $2^{\mathcal{O}(k\log k)}+n^{\mathcal{O}(1)}$ algorithm based on it and show that this FPT algorithm is asymptotically optimal under the ETH. Further, we generalize the problem by including an additional positive integer $c$ in the input that naturally captures how much agents can modify a given cycle packing to reject it. For every constant $c$, the resulting problem simplifies from being $\Sigma_{2}^{P}$-complete to NP-complete. The super-exponential lower bound already holds for $c=2$, though. We present an ad-hoc single-exponential algorithm for $c=1$. These results reveal an interesting discrepancy between the classical and parameterized complexity of the problem and give a good view of what makes it hard.
@@ 1 Introduction
For patients with severely reduced kidney function, transplantation is often the preferred treatment [7]. As it is possible to live with only one healthy kidney, live kidney donation is employed to reduce waiting times for sparsely available donor organs. In this setting, a willing donor offers one of their kidneys for transplantation to a patient. The patient and donor must be medically compatible for a successful transplantation. When the patient and donor are not compatible, they can opt to participate in a kidney exchange program (KEP).
Such programs aim to link multiple patient-donor pairs so that a donor from one pair is compatible with a patient from another pair. Their goal is to group sets of patient-donor pairs, such that within each group, a cyclic order of compatibility exists. Performing this cycle of transplants means that donors only undergo surgery if their corresponding patient receives a kidney in that exchange. Individual programs each have their own additional restrictions [5]. It is common to restrict the length of transplantation cycles to some constant $d$, due to practical considerations which include doing the associated surgeries simultaneously.
We study these KEPs from an algorithmic perspective. Given a set of participating patient-donor pairs and the medical compatibilities between them, the algorithmic task becomes to find a combination of transplantation cycles that maximizes the total number of patients receiving a healthy kidney. We can model each patient-donor pair as a vertex and the compatibility of the donor of a pair $a$ to the patient of a pair $b$ as a directed edge from $a$ to $b$. We call a cycle of length at most $d$ a $d$-cycle and use this notation to obtain the following algorithmic graph problem that describes KEPs in one of their most basic forms.
@@
This problem has already been studied from the perspective of parameterized algorithms in the literature, with fixed-parameter tractable (FPT) algorithms w.r.t. the input parameter $k$ and the number of neighborhood equivalence classes as examples [16, 20].
@@
Extensions of the standard kidney exchange problem have also been studied. In this work, we consider one such extension which we now motivate and introduce. It is known that the fraction of patients to receive a healthy kidney can increase significantly with an increase in the total number of participating pairs [2]. This motivates collaborations between multiple programs, from now on referred to as agents. Such collaborations, however, can be hindered by a mismatch between the individual interests of agents, as these need not align with the collective goal of helping as many patients as possible overall. In particular, a socially optimal solution—in which as many patients as possible are treated overall—is not guaranteed to help as many people from each individual agent as that agent could have achieved on their own. This may stop agents from participating in collaborative programs or, when they do participate, move them to withhold information from the collective.
Several ways of incorporating fairness into multi-agent KEPs have been studied [2, 3, 4, 8, 9, 15, 19]. Recently, Blom, Smeulders, and Spieksma [6] also introduced a framework to capture fairness via a game-theoretical concept of rejections. They translate this into an algorithmic task that asks for a solution that is not rejected by any agent. We study the resulting algorithmic task. To introduce it, we first define the concept of rejections.
@@
This definition captures that the rejecting agent $i$ could modify a given cycle packing $\mathcal{C}$ into another packing $\mathcal{C}^{\prime}$ that benefits more of their own patients by only making changes pertaining to patients within their control. We call a packing that is not rejected by any agent rejection-proof. Then, a guarantee to always find such a solution also serves as a guarantee for agents that participation in the joint program is never detrimental to their personal interests. This led Blom et al. [6] to introduce the following problem, whose parameterized complexity we investigate in this work.
@@
Rejection-Proof Kidney Exchange with $d$-cycles (RPKE-$d$) A directed graph $G$ where $V(G)=V_{1}\cup\ldots\cup V_{p}$ is partitioned among $p$ agents, and an integer $k$. Determine whether there is a rejection-proof set $\mathcal{C}$ of pairwise vertex-disjoint $d$-cycles such that $\mathcal{C}$ covers at least $k$ vertices. For ease of notation, we refer to any packing of pairwise vertex-disjoint $d$-cycles that covers at least $k$ vertices in an RPKE-$d$ instance as a candidate solution. Hence, a packing of $d$-cycles forms a solution to an instance if it is both a candidate solution and rejection-proof.
The above problem takes a leap in complexity compared to the single-agent problem KE-$d$. Whereas KE-$d$ is known to be NP-complete for all finite $d\geq 3$ [1], Blom et al. have shown RPKE-$d$ to be $\Sigma_{2}^{P}$-complete for $d\geq 3$, even when there are only two participating agents [6, Theorem 1]. In fact, it can be shown that checking whether a given agent rejects a given candidate solution is already NP-hard since this is at least as hard as checking whether a given packing is optimal in the NP-hard setting of only one agent. Partly motivated by this observation, we also consider a more restricted definition of rejections.
For some integer $c\geq 0$, we say that an agent $c$-rejects a packing $\mathcal{C}$ if the rejection criterion from 1.1 can be satisfied with $|\mathcal{C}_{\mathrm{rej}}|\leq c$. We call a packing that no agent $c$-rejects $c$-rejection-proof. Now, for each pair of positive integers $c$ and $d$ we define the problem $c$-Rejection-Proof Kidney Exchange with $d$-cycles ($c$-RPKE-$d$) as the variant of the above RPKE-$d$ problem in which we aim to determine whether the input contains a candidate solution that is $c$-rejection-proof.
@@
We study the complexity of RPKE-$d$ and $c$-RPKE-$d$ parameterized by $k$, the number of vertices to be covered, and we present both positive and negative results. We prove our positive results on related set packing problems that generalize the two graph problems. This allows us to use some of the tools developed for such problems. Prior to our work, the only positive result known for RPKE-$d$ was a polynomial-time algorithm for $d=2$ [8].
@@
We give an algorithm that, for constant values of $d$, solves RPKE-$d$ in $2^{\mathcal{O}(k\log k)}+n^{\mathcal{O}(1)}$ time on $n$-vertex graphs. It consists of two parts, the first of which is a kernelization algorithm. It takes an arbitrary instance as input and outputs an equivalent instance of size $\mathcal{O}\left((k^{d+1})^{d}\right)$ in polynomial time, via a careful agent-aware application of the sunflower lemma [12] to eliminate transplantation cycles. In set packing problems, removing sets during preprocessing typically comes at the risk of removing sets that are required to make an optimal solution, thereby possibly transforming YES-instances into NO-instances. The sunflower lemma provides a tool to circumvent this. Our rejection-proof problem settings introduce an additional type of behavior that needs care: transplantation cycles that do not belong to optimal solutions may still play a role in rejections. Removing the wrong cycles may cause an agent not to reject a certain candidate solution that it otherwise would have. We present a two-step reduction approach guaranteed to preserve the answer. The second part of the algorithm consists of brute-force on the reduced instance.
@@
We give a reduction from Subgraph Isomorphism to show that, under the Exponential Time Hypothesis (ETH), the algorithm is asymptotically optimal in terms of $n$ and $k$. We even show the stronger bound (independent of $k$) that a $2^{o(n\log n)}$ time algorithm for RPKE-$d$ or $c$-RPKE-$d$ would violate the ETH for any $c\geq 2$ and $d\geq 3$. We exploit the mechanism of rejections to facilitate our reduction from the Subgraph Isomorphism problem, for which such a superexponential ETH-based lower bound is known [10]. This gives the rejection-proof kidney exchange problems a similar complexity status to other problems in the literature, such as Metric Dimension and Geodetic Set parameterized by vertex cover number, for which a brute-force algorithm on a polynomial kernel is also ETH-optimal [14].
@@
To complement the ETH-hardness of $c$-RPKE-$d$ with $c\geq 2$, we give a $3^{n}\cdot n^{\mathcal{O}(1)}$ time algorithm for the problem when $c=1$. This allows us to see surprising jumps in the algorithmic complexity of $c$-RPKE-$d$ for varying values of $c$. While no big difference in complexity is obtained between $c=1$ and $c=0$ (in which case the problem can be solved in $2^{n}\cdot n^{\mathcal{O}(1)}$ time with a standard application of dynamic programming over subsets [10, Chapter 6]), our results show that the problem does become computationally harder from a parameterized perspective when $c=2$ (assuming the ETH). We also see that no more jumps in parameterized complexity appear for larger values of $c$. The surprising aspect here is that this one discrete jump in the parameterized complexity of $c$-RPKE-$d$ does not coincide with the jump in the classical complexity of the problem between being contained in NP vs. not being contained in NP: we have seen that, for constant $d\geq 3$, the $c$-RPKE-$d$ problem is contained in NP for every constant value of $c$, while the RPKE-$d$ problem is $\Sigma_{2}^{P}$-complete.
@@
The remainder of this work is organized as follows. We start with some preliminaries in section 2. Next, we give a polynomial kernel and an FPT algorithm for a set packing generalization of RPKE-$d$ in section 3, and we present a matching ETH-lower bound in section 4. In section 5, we give a single-exponential algorithm under $1$-rejections and we conclude in section 6. The proofs of statements marked $(\bigstar)$ are deferred to the appendix.
@@ 2 Preliminaries
All graphs considered are directed, unless stated otherwise. We use standard notation for graphs and concepts from parameterized complexity; see the textbook by Cygan et al. [11]. We write $\langle v_{1},\ldots,v_{\ell}\rangle$ to denote the directed cycle with edge set $\{(v_{1},v_{2}),\ldots,(v_{\ell-1},v_{\ell}),(v_{\ell},v_{1})\}$. We use the term packing to refer to a collection of sets that satisfies certain pairwise disjointness conditions, such as collections of vertex-disjoint cycles in a graph or collections of pairwise-disjoint subsets of a universe. Since the disjointness conditions differ depending on the context, we always specify which condition holds for the packing at hand.
@@
Next, we present the sunflower lemma by Erdős and Rado [12] and a related definition of sunflowers in set systems. In a set system $(U,\mathcal{S})$, a sunflower $\mathcal{F}$ with $z$ petals and core $Y$ is a collection of sets $S_{1},\ldots,S_{z}\in\mathcal{S}$ such that $S_{i}\cap S_{j}=Y$ for all $i\neq j$. We refer to a set $S_{i}\setminus Y$ as the petal corresponding to set $S_{i}$ and we require all petals to be non-empty. We state a variant of it as proven in the textbook by Flum and Grohe [13, page 212].
@@
When analyzing the running time of an algorithm whose input is a graph, we denote the number of vertices and edges in the input graph by $n$ and $m$, respectively. For algorithms operating on a set system, we use $n$ to denote the number of elements in the universe and $m$ for the total number of sets in the system.